OpenStudy (anonymous):

use the method of substitution to solve the system of equations {4x+y=2, x-y=3}

OpenStudy (nirmalnema):

x=y+3 you can get this by second equation now put this value in place of x in 1 eq.. u will get 4(y+3)+y=2 solving further 4y+12+y=2 5y=-10 y=-2 put this value any of the eq... u will get x = 1